In each of the 4 races, the democrats have 60% chance of winning.



Assuming that the races are independent of each other, what is the



probability that:



a) The Democrats will win 0 races, 1 race, 2 races, 3 races, or 4



races?



b) The Democrats will win at least 1 race.



c) The Democrats will win a majority of the races.

1
Expert's answer
2021-12-07T07:07:38-0500

Let "X=" the number of wins of the Democrats out of 4 races:

X"\sim Bin(n,p)"

Given "n=4, p=0.60, q=1-p=0.40."


a)


"P(X=0)=\dbinom{4}{0}(0.6)^0(0.4)^{4-0}=0.0256"

"P(X=1)=\dbinom{4}{1}(0.6)^1(0.4)^{4-1}=0.1536"

"P(X=2)=\dbinom{4}{2}(0.6)^2(0.4)^{4-2}=0.3456"

"P(X=3)=\dbinom{4}{3}(0.6)^3(0.4)^{4-3}=0.3456"

"P(X=4)=\dbinom{4}{4}(0.6)^4(0.4)^{4-4}=0.1296"

b)


"P(X\geq 1)=1-P(X=0)"

"=1-0.0256=0.9744"

c)


"P(X>2)=P(X=3)+P(X=4)"

"=0.3456+0.1296=0.4752"